OpenStudy (anonymous):

the manager of a coffee shop has one type of coffee that sells for $5 perlb and another type that sells for $14 perlb.the manager wishes to mix 40 pounds of the $14 coffee to get a mixture that will sell for $8 perlb.how many pounds of the $5 coffee should be used?

OpenStudy (amistre64):

b@5$ = 40(14-8)/(5-8) maybe? gotta work out the bugs if not :)

OpenStudy (amistre64):

40(14) + b(5) = (40+b)(8) 40(14) + b5 = 40(8) +b8 40(14-8) = b(8-5) b = 40(14-8)/(8-5) ... its good, just woulda gotten a negative

OpenStudy (amistre64):

so what do we get?

OpenStudy (anonymous):

I don't have an answer

OpenStudy (anonymous):

how many pounds of the $5 coffee would be used?

OpenStudy (amistre64):

well, i gave you the formulations and filled in the numbers ... i even calculated the results to check with. The only thing I havent done for you in this is just blurt out the answer :)